The sum of 8 and a certain number is equal to the product of the number and 3.find the number
Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52777)

.

    8 + x = 3x

      8   = 3x - x

      8   =    2x

      x   =    8/2 = 4.


ANSWER.  The number is 4.

Solved.



Answer by math_tutor2020(3816)


x = unknown number
x+8 = sum of the number and 8
3x = product of the number and 3

Those two mentioned expressions are equal to one another
x+8 = 3x
8 = 3x-x
8 = 2x
2x = 8
x = 8/2
x = 4 is the unknown number

In the second step, I subtracted x from both sides. A bit afterward, I divided both sides by 2 so I can isolate x.

Check:
x+8 = 4+8 = 12
3x = 3*4 = 12
Both sides get us the same value, so we've shown that x = 4 is the solution to x+8 = 3x

Answer: 4